### v1

# Does Time Have Its Place? Temporal Heads: Where Language Models Recall Time-specific Information

Abstract
While the ability of language models to elicit facts has been widely investigated, how they handle temporally changing facts remains underexplored. We discover Temporal Heads, specific attention heads primarily responsible for processing temporal knowledge through circuit analysis. We confirm that these heads are present across multiple models, though their specific locations may vary, and their responses differ depending on the type of knowledge and its corresponding years. Disabling these heads degrades the model’s ability to recall time-specific knowledge while maintaining its general capabilities without compromising time-invariant and question-answering performances. Moreover, the heads are activated not only numeric conditions (“In 2004”) but also textual aliases (“In the year …"), indicating that they encode a temporal dimension beyond simple numerical representation. Furthermore, we expand the potential of our findings by demonstrating how temporal knowledge can be edited by adjusting the values of these heads [^1] Our datasets and code are publicly available at https://github.com/dmis-lab/TemporalHead.
Does Time Have Its Place?
Temporal Heads: Where Language Models Recall Time-specific Information
Yein Park1, Chanwoong Yoon1, Jungwoo Park1,3, Minbyul Jeong2thanks: Corresponding authors, Jaewoo Kang1,311footnotemark: 1 Korea University1  Upstage AI2  AIGEN Sciences3 {522yein, cwyoon99, jungwoo-park, kangj}@korea.ac.kr
1 Introduction
“Remembrance of things past is not necessarily the remembrance of things as they were.” (Proust, 1992)
This profound and intricate relationship between memory and truth resonates deeply with one of the central challenges in modern artificial intelligence. While large language models (LLMs) like GPTs (OpenAI, 2022, 2024a, 2024b) and LLaMA familiess (Touvron et al., 2023a, b; Dubey et al., 2024) have demonstrated remarkable capabilities in leveraging factual knowledge, they face a unique challenge that mirrors human memory: the accurate representation of temporal knowledge—facts that transform across different time points.
Unlike static facts (e.g., “The capital of France is Paris”), many real-world facts change over time (e.g., a politician’s term in office, a sports player’s team membership in a given year). This time-evolving nature necessitates that LLMs accurately capture such change. To do so, they must not only track newly updated facts within a specific timeline, but also retain historical information across different time periods (Jang et al., 2022). This presents a significant challenge, as models must contend with tracking and reasoning over temporal changes in knowledge (Kasai et al., 2023). However, beyond prompting (Mitchell et al., 2022; Park et al., 2025) or retrieval-augmentated generation (Lewis et al., 2020; Gutierrez et al., 2024), the internal mechanisms by which models adapt to temporally evolving facts remain relatively underexplored.
Figure 1: Temporal Heads exist within various TKCs at different times $T_{k}$. Ablating them disrupts the model’s temporal alignment, yielding incorrect objects.
Figure 2: Overview of temporal knowledge circuit analysis. (A): Construct temporal knowledge circuits (TKCs), and compare it with general knowledge circuits (KCs) using time-invariant knowledge. Circuits reproduce residual streams for time T, subject S and relation R. This verifies temporal heads only found in each different TKCs of various year $T_{k}$. (B): Example of simplified TKC. Here, basic knowledge nodes is colored violet, (common in both), while Temporal Heads is highlited. (C): Attention map for temporal heads. a15.h0 means the 15th layer’s first attention head. Each head’s attention pattern is represented as the output logits of the hean by mapping to vocabulary space. Queries are input tokens focusing on others, while keys are the tokens being focused on. Values represent attention weights, indicating the strength of this focus. Total results are in Figures 78 and 911.
Empirical observations suggest that LLMs already possess some level of temporal awareness (Nylund et al., 2023; Mousavi et al., 2025). This raises the question of whether the model is inherently capable of encoding and utilizing temporal knowledge. For instance, when prompted with time-specific queries like “In 1999, [X] was a member of sports team”, the model may generate the correct team [Y] relevant to that year, indicating that certain time-conditional links are embedded in its internal parameters. The key puzzle, however, is how this temporal knowledge is organized and recalled. Do LLMs internally have a place for Time, adjusting their factual outputs based on the input time condition? If so, where within the model architecture—among the attention heads and feed-forward layers—does this mechanism reside?
To address them, we apply Circuit Analysis (Elhage et al., 2021; Wang et al., 2023) to reconstruct the model’s computations via localized subgraphs of attention heads, feed-forward networks (FFN), and residual streams. Especially, by systematic ablating (zeroing out) attention heads or MLP components, it pinpoints which parts are responsible for eliciting knowledge in each knowledge recalling inference (Yao et al., 2024). These knowledge circuits enable to measure how much each nodes or edges in subgraph contribute to processing facts.
We extend it into temporal dimension, capturing how models reacts to time-evolving attributes using Temporal Knowledge Circuits (Figure 1 (A)). We then identify Temporal Heads, such as a15.h0 and a18.h3, which are exclusively activated for temporal knowledge while remaining inactive for time-invariant information. Each model have its own temporal heads, which exhibit a strong influence on temporal input tokens in attention maps. Moreover, ablating these heads significantly reduces time-specific factual accuracy, leading to temporal mismatches as suggested in Figure 1 (B).
One step further, we explore in-depth impacts of temporal heads among different years, knowledge and conditioning types. Ablating them exclusively affects temporal information, while having negligible impact on time-invariant knowledge and general QA performance. Notably, these temporal heads are activated for both numerical expressions (“In 2004”) and textual conditions (“In the year the Summer Olympics were held in Athens”), indicating that they encode a broader temporal dimension beyond simple numerical representation. Building on this, we demonstrate that temporal knowledge editing-selectively adding their activations-enables direct intervention in year-conditioned factual recall. Through this targeted manipulation, our experiments demonstrate that the temporal heads serve as key subcomponents for encoding and modifying time-sensitive knowledge.
2 Preliminaries
In this section, we provide background on the Circuit Analysis (Olah et al., 2020; Nanda et al., 2023; Conmy et al., 2023), which represents the model’s computation through structured subgraph of its components.
2.1 Circuit Analysis
Circuit analysis represents a transformer’s computation as a directed acyclic graph (DAG) $G=(N,E)$, where each node in $N$ corresponds to a distinct component in the model: attention heads $A_{l,j}$ (at layer $l$ and head $j$), MLP modules $M_{l}$ for each layer, the input node $I$ (embeddings), and the output node $O$ (logits). Thus, we formally define the set of nodes as:
$N=\{I,A_{l,j},M_{l},O\}.$ (1)
The edges in $E$ represent residual connections that propagate activations between these nodes:
$E=\{(n_{x},n_{y})\mid n_{x},n_{y}\in N\}.$ (2)
A circuit is defined as a subgraph $C\subseteq(N,E)$ selected to explain a specific behavior of interest–for instance, how certain tokens influence the model’s output or how factual knowledge is stored and elicited. By examining which nodes and edges are crucial for producing a particular prediction, we can identify the subgraph (the circuit) that governs each behavior.
2.2 Knowledge Circuit
A knowledge circuit (Yao et al., 2024) focuses on how a model treats the subject $s$, and relation $r$ to generate the object $o$ using a knowledge triplet $(s,r,o)$. By systematically ablating (i.e. zeroing) parts of the model, it identifies the crucial nodes responsible for this generation and constructs a subgraph $KC\subseteq(N,E)$ whose removal breaks the model’s ability to produce the correct object. Concretely, it define a performance metric as:
$\begin{split}S(e_{i})=&\;\log\bigl{(}p_{G}(o\mid s,r)\bigr{)}\\ &-\log\bigl{(}p_{G/e_{i}}(o\mid s,r)\bigr{)}.\end{split}$ (3)
where $p_{G/e_{i}}$ denotes the model’s probability of next-token prediction after ablating (i.e. zeroing) the activation of a node or edge $e_{i}$. If $S(e_{i})$ exceeds a threshold $\tau$, $e_{i}$ is deemed critical and retained in $KC$; otherwise, $e_{i}$ is pruned. This yields a minimal set of heads/MLPs whose connections critically shape the binding of $(s,r)$ to the correct answer $o$.
Unlike a generic circuit for any functionality, a knowledge circuit specifically captures the local subgraph dedicated to storing and relaying factual content for the knowledge triplet at hand. We specifically utilize effective attribution pruning-integrated gradients (EAP-IG), which ablating (zeroing) candidate edges and measuring drops in correct prediction (Hanna et al., 2024). For more details, see Appendix 8.1.
3 Knowledge Circuit Deciphers Temporal Head in LLMs
We now explore how knowledge circuits, extracted via EAP-IG pruning, can reveal specialized Temporal Heads in large language models (LLMs). We extend knowledge circuits in 2.2 to temporal knowledge circuits by analyzing how the same subject–relation pair can produce different objects across multiple time points. Specifically, we seek to identify which edges encode time-dependent specificity, such that an edge $e_{i}$ is crucial for predicting the time-relevant object $o_{k}$ at period $T_{k}$. Given a knowledge circuit score $S(e_{i})$ (Eq. 3), we define its temporal variant as follows:
$\begin{split}S(e_{i},T_{k})=&\;\log p_{G}(o_{k}\mid s,r,T_{k})\\ &-\log p_{G/e_{i}}(o_{k}\mid s,r,T_{k})>\tau.\end{split}$ (4)
where $T_{k}$ indicates a specific time (or period), and $o_{k}$ is the corresponding object for subject $s$ and relation $r$ at time $T_{k}$. Thus, $S(e_{i},T_{k})$ measures the contribution of edge $e_{i}$ to correctly predicting $o_{k}$ under time $T_{k}$. For highlighting importance and simplifying graphs, edges retained in the temporal circuit satisfy $S(e_{i},T_{k})>\tau$, ensuring they encode time-dependent knowledge. Here, we decide to attach temporal conditioning in front of subject, following prior insight from causal tracing (§8.2) and details in Appendix 8.3.
3.1 Implementations
We conduct experiments primarily on three LLMs: Llama-2-7b-chat-hf (Touvron et al., 2023b), Qwen1.5-7B-Chat (Bai et al., 2023; Team, 2024), Phi-3-mini-4k-instruct (Abdin et al., 2024). We adopt transformer lens (Nanda and Bloom, 2022) to intercept and ablate model components, enabling EAP-IG-based circuit discovery. We mainly illustrate results on Llama2, though similar trends emerge in the other models. More details are described in Appendix 8.1.1.
3.1.1 Circuit Reproduction Score
To evaluate how well a pruned circuit reproduces the full model’s behavior, we define the Circuit Reproduction Score (CRS), ranging from 00 to $100$. Let $B$ be the baseline performance of the full model on time-conditioned prompts, and $P$ be the performance of the pruned circuit. If the pruned circuit maintains or exceeds the baseline performance ($P\geq B$ when $B>0$), we assign it the maximum CRS as follows:
$\mathrm{CRS}(B,P)=100.$ (5)
Otherwise, the score follows an exponential decay:
$\mathrm{CRS}(B,P)=100\times\sigma\exp\left(-\alpha\frac{d}{|B|}\right),$ (6)
where $d=\max\{B,0\}$. The factor $\sigma\in(0,1]$ accounts for sign mismatches, adjusting for cases where the pruned circuit’s output deviates in direction from the full model. A higher CRS indicates better reproduction of the full model’s predictions. Details on hyperparameters and adjustments are deferred to the Appendix 8.4.
3.2 Dataset
Our dataset comprises (statistics in Appendix 8.5):
  • Temporal Knowledge: Various categories of knowledge samples that embed a specific year (e.g., 1999, 2004, and 2009) alongside a factual statement (e.g., which sports team or president is correct in that year) based on Wikidata (Vrandečić and Krötzsch, 2014).
  • Time-Invariant Knowledge: Commonsense data from LRE (Hernandez et al., 2024) (e.g., object superclass, fruit inside color), plus newly implemented numerical facts embedded in subject/object (e.g., geometric shape or roman numerals). These tasks assume no explicit time-based shift.
  • Unstructured QA: We utilize TriviaQA (Joshi et al., 2017) and Math (Wang, 2022) QA in ChroKnowledge (Park et al., 2025) for unstructured, general QA to verify the ablation effect with basic LLM’s tasks.
For each data point, we run both a clean prompt and a corrupted prompt, following EAP-IG guidelines. We focus on the first token(s) that differ, capturing the key transition that determines correctness. In the QA setting, we evaluate models using standard TriviaQA validation metrics, including exact match (EM) and F1 scores. For Math ChroKnowledge, we employ a multiple-choice QA (MCQA) template, scoring responses based on probability (%). Given that models possess some degree of inherent knowledge (Yao et al., 2024), we assess their performance under zero-shot and greedy decoding.
Category Knowledge #Node #Edge CRS
Temporal
Sports Nicolas Anelka 29 37 74.14
David Beckham 43 80 39.53
Presidents Argentina 42 102 60.97
South Korea 46 110 65.55
CEO Hewlett-Packard 52 115 53.49
Chrysler 51 97 57.10
Defense United States 50 137 48.08
China 19 19 37.62
Avg 42 87 54.56
Time-Invariant
CommonSense Object Superclass 43 56 44.47
Conditional CS Fruit Inside Color 76 131 53.08
Num in Obj Geometric Shape 52 118 76.09
Num in Sub Roman Numerals 43 135 95.70
Avg 54 110 67.33
Table 1: Statistics of temporal knowledge circuits for Llama2, both temporal and time-invariant knowledge. For temporal knowledge, each type of knowledge is reproduced with three selected years: 1999, 2004, and 2009. The numbers of nodes, edges and CRS is the average of each knowledge’s yearly circuits.
3.3 Evaluation
After pruning less-contributory nodes via EAP-IG, we measure how well the resulting subgraph preserves the model’s original performance on each knowledge type. Table 1 and 34 shows the average number of nodes and edges in these pruned circuits, along with their CRS. We then apply threshold $\tau$ to remove edges/nodes that contribute marginally to object prediction, retaining only edges with scores above $\tau$ and their corresponding nodes.
In Llama2, both temporal and time-invariant knowledge circuits effectively capture the model’s internal knowledge flow, with average CRS exceeding 50 in both cases. However, temporal circuits exhibit more variability, likely due to the inherent complexity of year-based facts. These tasks demand precise temporal conditioning, adding an extra difficulty, not just simply generating any possible objects. Even when models are expected to retain such knowledge, the increased complexity underscores the nuanced nature of temporal reasoning compared to time-invariant knowledge.
3.4 Findings
We now identify common nodes in all circuits (e.g., [input], [logits], MLP m2, m24, m30, etc.) and a set of temporal-only nodes that appear exclusively in circuits for year-dependent prompts as in Figure 2. Firstly, most MLP nodes were appeared both temporal and time-invariant knowledge, as they are activated for storing knowledge (Geva et al., 2021; Dai et al., 2022; Niu et al., 2024).
What is impressive stood out in the attention heads. Temporal Heads, appearing in almost every temporal knowledge circuits but not time invariants, are shown: a15.h0, a18.h3 in Llama2. Those temporal heads reoccur across multiple year-specific circuits, and it is different for other model’s cases like a17.h15 for Qwen 1.5 in Table 2. Visualizing their attention maps in Figure 2 (C) indicates a strong focus on “In 19xx” and subsequent subject phrases, as key tokens revolve around temporal conditions with queries hooking into the subjects. This pattern corroborates the idea that these heads facilitate year-subject binding—justifying the label “temporal”, as this kind of task specific attention heads were previously suggested by Wang et al., 2023; Merullo et al., 2024; Chughtai et al., 2024; Wu et al., 2024 and  Zheng et al., 2024.
When lowering the ratio of exhibition (e.g., 70-80%), additional heads (e.g., a0.h15, a20.h17, a31.h25) emerge. These Backup Temporal Heads are also exclusive to temporal knowledge circuits, though their emerging varies different among types of knowledge and years. But interestingly, even at high ratio, no heads are exclusive in time-invariant knowledge circuits. This suggests that many “general knowledge” heads overlap with or are reused by knowledge recalling tasks, whereas certain specialized heads exist only for time-based tasks.
Figure 3: Log probability results with temporal knowledge; In XXXX, the president of South Korea was. (A) shows prediction probability change among results of Llama2. The effect of head ablation reacts differently for each selected year with the same prompt. Each subplot in (A) represents the probability distribution of correct (green) and incorrect (red) predictions, where the x-axis denotes probability values and the y-axis differentiates between target and non-target responses. Total results for each model are in Figures 1213 in Appendix. (B) illustrates the performance degradation trends across various years. As averaging the result of ablation, the gray space between two line plots represent degradation level pointed out by red arrows (which becomes darker and bigger when the gap is wider). The background shows how objects were changed in the time range between 1999 to 2009.
In the next (§4), we delve into further ablation experiments to verify that ablating temporal heads indeed degrades year-specific predictions, reinforcing their role as the crucial channel through which the model recall knowledge conditioned on time.
4 In-Depth Analysis of Temporal Heads
We conduct a more fine-grained analysis to understand how temporal heads identified in the extracted circuits impact final predictions, especially for temporally changing facts. Drawing inspiration from Borchmann 2024 on log-probability based evaluation, we perform targeted Attention Head Ablation Inference4.1) to observe how the model’s confidence shifts when certain “temporal” heads are zeroed out. We then test an Alias scenario with temporal conditioning in textual context (§4.2) to see if the same heads reappear for less explicit time references. Finally, we explore a Temporal Knowledge Editing5) that uses attention addition to reinforce or awake year-specific content.
THs Settings Temporal (%) Invariant (%) QA (F1)
Llama-2-7b-chat-hf
a18.h3, a15.h0 Baseline 29.7 61.8 55.4
Ablation 25.6 $\Downarrow$ 61.7 $\downarrow$ 54.9 $\downarrow$
Qwen1.5-7B-Chat
a17.h15 Baseline 22.4 62.7 49.7
Ablation 19.8 $\Downarrow$ 62.6 $\downarrow$ 49.5 $\downarrow$
Phi-3-mini-4k-instruct
a10.h13 Baseline 35.4 59.8 46.8
Ablation 26.0 $\Downarrow$ 60.6 $\Uparrow$ 46.2
Table 2: Temporal Heads (THs) across different LLMs. The scores besides each heads are evaluated in three cases (temporal knowledge, time-invariant knowledge, and TriviaQA) with two settings (baseline inference and ablation inference). Scores are checked with the average performance for each tasks, measured in probability (%) or f1 score. While performance in temporal knowledge drops significantly (3 to 9%), time-invariant and general QA remain relatively stable or even goes up.
4.1 Attention Head Ablation Inference
Motivation
While temporal knowledge circuit construction based on EAP-IG pruning (§3) reveals the structure of temporal knowledge processing, we still need direct evidence that certain “temporal heads” genuinely mediate year-based predictions. We adopt a hard-coded approach that sets the selected attention head’s output weights to zero, thus preventing it from contributing to the residual stream. We then measure changes in the model’s log probability for the correct target object vs. competing objects in different time.
Log Probability Variation
Following Borchmann 2024, we assess temporal knowledge retention by evaluating changes in object probabilities under head ablation. Let ${O}$ be the set of all candidate objects (e.g., teams, presidents) in the time range, and $p(o|s,r,T)$ the model’s probability of selecting object $o$ from subject $s$, relation $r$ and time $T$. The model’s default choice is labeled Target if it matches the correct temporal fact, otherwise Non-Target. After ablating suspected temporal head(s), we recompute object probabilities:
$z_{o}$ $=\log p_{\text{ablate}}(o|s,r,T),$ (7)
$\hat{p}_{o}$ $=\frac{\exp(z_{o})}{\sum_{o^{\prime}\in O}\exp(z_{o^{\prime}})},$ (8)
where $p_{\text{ablate}}$ denotes the log-probability computed by forward pass of model, ablating corresponding heads. This evaluates how the probability distribution over $O$ shifts, rather than just predicting the most likely answer. Details in Appendix 8.6.
4.1.1 Result of Temporal Knowledge
As shown in Figure 3 (A), ablation significantly reduces log probability for the correct year-specific Target in temporal tasks. When ablating a15.h0 or a18.h3 or both of them, the model frequently chooses Non-Target objects from ${O}$ (e.g., a president of different year). Not just raising of those percentage, specific attention heads influence each years differently; some are more critical for 1999, while others have a stronger effect in 2004 or 2009. For instance, ablating a18.h3 significantly impacts 2004 but has a lesser effect on 2002.
Figure 3 (B) illustrates the varying degrees of performance degradation across different years. The red arrows highlight these degradation levels, where darker and thicker arrows indicate a more pronounced effect of ablation. Notably, around object transition periods (e.g., between 2002–2003 and 2007–2008), the non-target probability spikes, confusing when knowledge boundaries shift along the timeline. This aligns with the intuition that temporal knowledge transitions introduce uncertainty in the model’s predictions in temporal context.
4.1.2 Result of Time Invariant Knowledge
By contrast, ablating the same heads for invariant knowledge (e.g., fruit inside color) causes minimal performance drop in Table 2 and Figure 4. This indicates that “temporal heads” indeed route only temporally conditioned knowledge, and disabling them forces the model to make temporally incorrect rather than incorrect of stable knowledge. Besides, Phi-3-mini affects more sensitively than others as its parameter size is half of other two models, resulting more reactive to small changes in attention alignment. This even causes a slight gain of performance in time-invariant knowledge tasks.
4.1.3 Result of General QA
As Table 2 and result in Appendix 8.7 shows, ablating temporal heads doesn’t harm common knowledge recalling or answering general knowledge questions. Here, we test TriviaQA and Math ChroKnowledge and find out that just ablating temporal heads doesn’t effect the performance of basic QA, droping almost less than 0.6 in f1 score.
Figure 4: Head ablation effect across various knowledge types. Three selcted model shows distinct differentiation for temporal knowledge (left side) and time invariant knowledge (right side). The change of performance is calculated with the average score of baseline (non-ablation) and modified (ablated result), using model specific temporal head information. While degrees of degradation is different among models, overall tendency reflects the importance of temporal head to inference temporal knowledge.
4.2 Alias Test With Textual Conditioning
In previous findings of Section §3.4, we experimented with cases where numeric values were present either in the prompt (Roman Numerals) or in the answer object (Geometric Shape) under time-invariant conditions (like “Triangle has 3 sides”). For all scenarios, temporal heads did not emerge, suggesting that their activation is not merely a response to numerical information but rather specific to temporal knowledge processing. We further investigate whether these same heads appear for less direct numeric conditioning. Instead of a literal “In 2004” prompt, we use “In the year the Summer Olympics were held in Athens” or “For his first,” providing an indirect textual condition referencing the relevant time. We again construct knowledge circuits and observe which heads surpass threshold.
Such “alias” statements yield smaller CRS (e.g., 40.3 in president cases), though, temporal heads still appears. These heads may not always exceed normal threshold (e.g. $\tau=0.1$), they still register moderate importance. Coupled with results from the numeric “In 2004” prompt, this indicates that those heads do not rely solely on numeric tokens, but also respond—albeit less strongly—to textual or event-based temporal conditioning. This further validates that they encode a temporal dimension, rather than merely responding to arbitrary numbers. Visualized results are in Figure 14 of Appendix.
5 Temporal Knowledge Editing
Lastly, we explore an approach to confirm that injecting or amplifying temporal head’s attention value can effectively “edit” year-specific knowledge as in Figure 5. Given a source_prompt (where the model is confident about a certain year’s fact) and a target_prompt (where it confuses the same year) based on log probability results, we:
  1. 1.
    Extract the value of attention head $\mathbf{a}_{\mathrm{src}}$ from the source_prompt at a chosen layer/head (e.g. a18.h3).
  2. 2.
    Average over total source prompts (e.g., "In 2009, the name of president of South Korea was").
  3. 3.
    Inject the modified attention value into the target_prompt at the corresponding temporal token position, scaled by a coefficient $\lambda$:
Details of adding an attention is in Appendix 8.8.
Figure 5: Example Of Temporal Knowledge Editing. From the source prompt, we catch the specific attention value of model’s head, for example, a18.h3. By simply adding it to target prompt, the model’s output is changed into temporally correct answer from temporally wrong answer. The headmap below denotes the number of success in editing for every combination of layers and heads. The most successful case in here is temporal heads a18.h3 as highlighted, following other heads such as backup temporal heads a20.h17.
This modification is applied dynamically using a forward hook mechanism at inference time, preserving the overall model parameters while selectively influencing time-conditioned factual recall. We test it with model wrong answer in a normal condition, varying the injection coefficient across three cases ($\lambda=1,3,6$), following Turner et al., 2023; Rimsky et al., 2024, which emphasized its impact.
Remarkably, the model’s completions shift from a temporally incorrect response (“changed to Vladimir Putin”) to the correct one (“Dmitry Medvedev”), aligning with the known presidency timeline. The heatmap in Figure 5 further supports this by visually representing the effectiveness of temporal knowledge editing across all layers and heads. While certain attention heads can influence the model’s response, the most successful cases are consistently linked to temporal heads, with a18.h3 exhibiting the highest success rate. Additionally, backup temporal heads, such as a20.h17, also rank among the top-performing heads, reinforcing their critical role in preserving and modifying time-conditioned knowledge. This highlights that temporal factual recall is not arbitrarily distributed but is instead concentrated in specialized subcomponents. Other results are in Figure 15.
This targeted intervention remains minimally invasive, as it does not require global fine-tuning but instead modulates the value of a single specialized head, thereby preserving most of the model’s prior knowledge. Taken together, these findings reinforce the hypothesis that LLMs harbor a temporal subcomponent within specialized attention heads. By intercepting or amplifying these temporal heads, we can selectively alter time-conditioned responses, strengthening the claim that these heads are integral to the reinforcement of year-based factual knowledge.
6 Related Works
6.1 Temporal Knowledge of LLM
Despite advancements in LLMs, handling temporal knowledge remains a key challenge. While prior works focus on factual consistency (Petroni et al., 2019; Kassner and Schütze, 2020) or refining model editing in MLP layer (Mitchell et al., ; Meng et al., 2022, ), few address how facts evolve over time. Studies on time-aware QA and temporal probing (Chen et al., 2021; Zhang and Choi, 2021; Dhingra et al., 2022; Jang et al., 2022) reveal that LLMs struggle with dynamically shifting facts. Recent approaches attempt explicit temporal alignment (Kim et al., 2024; Zhao et al., 2024; Mousavi et al., 2025; Park et al., 2025), but have focused on external evaluations. Our findings highlight that LLMs encode temporal facts implicitly, relying on manipulable attention heads, underscoring the need for better temporal supervision and disentangled knowledge representations.
6.2 Attention Heads in Language Models
Under mechanistic interpretability Olah et al. (2020); Vig et al. (2020); Sharkey et al. (2025), researches about attention heads were done by Voita et al., 2019; Wang et al., 2023; McDougall et al., 2024, showing off specific heads that copy key tokens to the output, ensuring consistency in transformers. These Mover Heads are a kind of induction heads (Olsson et al., 2022) moving syntactic information (Ferrando and Voita, 2024). Other works were followed as finding out retreval heads (Wu et al., 2024), heads for semantic information for color (Merullo et al., 2024), or subject and relation (Chughtai et al., 2024). Those various kinds of attention heads attend to critical tokens and directly influence the logits by writing their embeddings into the residual stream (Zheng et al., 2024).
Experiments show that ablating those heads significantly disrupts tasks like syntactic induction or semantic information understanding, highlighting their specific roles. A special case, Backup Heads, remains inactive under normal conditions but replicates task specific head functionality when primary heads are ablated. This ensures model robustness by maintaining token copying behavior even when key circuit components are disrupted. We treat founded temporal attention heads as a subcategory of semantic heads like subject heads and relation heads (Chughtai et al., 2024) in our experiments.
7 Conclusion
We systematically investigate how LLMs can handle temporal knowledge, focusing on time-dependent facts. Through our experiments, we uncovered Temporal Heads that selectively mediate the activation of time-variant knowledge. Ablating these heads leads to temporal mismatches while leaving time-invariant knowledge and general QA performance unaffected. Note that these heads are also activated under textual conditioning, and using their value for editing successfully changes the models’ responses with minimal intervention.
As a foundational step, our work explores how LLMs can actively manage temporal information rather than merely integrating temporal context. We believe our analysis offers valuable insights into the inner mechanisms of LLMs and can inspire future approaches for time-aware model alignment and precise temporal updates by selectively targeting temporal heads, rather than relying on global retraining.
Limitations
While our approach demonstrates promising results in identifying and analyzing temporal knowledge circuits, we acknowledge some limitations in our current work. First, analysis of unstructured temporal QAs like General ChroKnowledge (Park et al., 2025) were constrained, as the underlying multiple-choice options in those tasks typically do not exhibit temporal dependencies. So we focused more on our temporal knowledge dataset, abundantly describing the effect of ablation in these cases.
On the other side, as EAP-IG didn’t support models with Grouped-Query Attention (GQA), which cannot use the split_qkv_input option, our main analysis exclude those models like Llama-3-8B-Instruct (Dubey et al., 2024). Still, we checked their results and found that even their CRS is not quite enough and their circuit construction is not detailed, temporal heads are still could be founded: a18.h15 and a23.h26.
Acknowledgments
This work was supported in part by the National Research Foundation of Korea [NRF-2023R1A2C3004176, RS-2023-00262002], the Ministry of Health & Welfare, Republic of Korea [HR20C002103], and the ICT Creative Consilience program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the MSIT [IITP-2025-2020-0-01819].